How many real number solutions does the equation have?y=3x^2-5x-5
Answer by rothauserc(4718)

0 = 3x^2-5x-5
the quadratic formula gives us
x = (5+square root(85)) / 6 = 2.369924076
x = (5-square root(85)) / 6 = −0.70325741
there are two real number solutions to the equation
***************************************************
substitute each solution in the original equation
(3×(2.369924076)^2)−(5×2.369924076)−5 = −0.000000002
(3×(−0.70325741)^2)−(5×−0.70325741)−5 = 0.000000004
our answers check